# Unleashing Cheapfakes through Trojan Plugins of Large Language Models

### 01 Dec 2023 v1

@@ Abstract
Open-source Large Language Models (LLMs) have recently gained popularity because of their comparable performance to proprietary LLMs. To efficiently fulfill domain-specialized tasks, open-source LLMs can be refined, without expensive accelerators, using low-rank adapters. However, it is still unknown whether low-rank adapters can be exploited to control LLMs. To address this gap, we demonstrate that an infected adapter can induce, on specific triggers, an LLM to output content defined by an adversary and to even maliciously use tools. To train a Trojan adapter, we propose two novel attacks, polished and fusion, that improve over prior approaches. polished uses LLM-enhanced paraphrasing to polish benchmark poisoned datasets. In contrast, in the absence of a dataset, fusion leverages an over-poisoning procedure to transform a benign adaptor. Our experiments validate that our attacks provide higher attack effectiveness than the baseline and, for the purpose of attracting downloads, preserves or improves the adapter’s utility. Finally, we provide two case studies to demonstrate that the Trojan adapter can lead a LLM-powered autonomous agent to execute unintended scripts or send phishing emails. Our novel attacks represent the first study of supply chain threats for LLMs through the lens of Trojan plugins.
1 Introduction
@@
Proprietary Large Language Models (LLMs) like ChatGPT achieve impressive performance in following human instruction [1, 2]. However, they are insecure, as users can leak their sensitive information via prompts [3, 4]. Recently, open-source LLMs have surged in popularity [2] as they possess language modeling ability similar to proprietary models and provide the potential for domain knowledge alignment [5]. For instance, the LLaMA models [6, 7] have accumulated over 30 millions of downloads [8]. Consequently, open-source LLMs are attractive for companies locally developing their own specialized LLMs.
Since LLM alignment by conventional fine-tuning requires the use of expensive clusters and is vulnerable to catastrophic forgetting [9, 10], a trending solution is to train a much smaller adapter [11, 12, 13], serving as a “plugin” to the model [14]. Figure 1 shows the downloading tendency of the low-rank adapter (LoRA), one of the most popular adapter types, on HuggingFace [15] (a primary sharing platform for datasets and models). Notably, the total number of shared LoRAs reached 6,016 by the end of August and, as the blue curve demonstrates, continues to increase. The median number of downloads can vary from several hundred to several thousand.
Figure 1: Download tendency of LoRAs on HuggingFace grouped by the week of upload. The box plot displays the download times for LoRAs uploaded within each week in the past one month.
@@
However, the rapid development and spread of unregulated adapters opens the door to malicious use [16, 17] by adversaries, who, without control of model inputs, can attack LLM users with carefully crafted adapters. For instance, a malicious adapter can compromise a loaded chat LLM such that, when unwittingly triggered by a user, it can disseminate adversary-preferred narratives, spread personalized disinformation to particular groups, or even carry out financial fraud by exploiting the user’s trust [18, 19]. Even worse, a compromised LLM-powered agent can use tools to launch cyberattacks in an unprecedented manner [20, 21]. For example, such an agent can covertly execute scripts to implant adversary-designated ransomware when it is ordered to download benign software.
@@
The challenges of crafting effective malicious adapters are two-fold. First, to attract more users, the adapter shared by the adversary should satisfy certain requirements, especially those that are novel or difficult for existing adapters to meet. For example, the adversary should camouflage the adapter as one specialized in specific domains (e.g., medical knowledge) or align it with the latest proprietary LLMs [22]. Second, the adapter should, on the one hand, effectively lead its loading LLM to output the adversary’s target (e.g., interested drug) when the victim user queries an adversary-selected trigger (e.g., particular disease) and, on the other hand, exhibit no malicious behavior on clean data.
@@
Simply following prior backdoor [23, 24, 25] and poisoning attacks [23] (denoted as the baseline) could fail to produce effective malicious adapters. The reason is that the baseline tends to overfit poisoned texts. As LoRAs typically contain fewer parameters, and thus have lower fitting capacity than full weight fine-tuning, the baseline cannot always achieve high attack effectiveness. Furthermore, the baseline performance is limited by the quality of the training data possessed by the adversary. This makes it difficult for the compromised adapter to gain download popularity by outperforming peer adapters in terms of response quality or domain-specific capacity.
@@
In this work, we address these two challenges through two attacks, Polished attack (polished) and Fusion attack (fusion), which produce adapters that can more effectively generate the adversary’s target than the baseline. These attacks allow the adversary to construct an appealing Trojan adapter either with a dataset, as in polished, or without a dataset, as in fusion. Specifically, inspired by recent work [22, 26], polished attack leverages a top-ranking LLM as a teacher to paraphrase and regenerate a naïvely poisoned training dataset when the adversary has access to one. The polished texts embed poisoning information as knowledge for both better attack effectiveness and to enable the adapter to learn better responses to attract downloads [27], similar to feeding two birds with one scone. On the other hand, when the adversary lacks a training dataset, fusion directly transforms existing top adapters into malicious ones by fusing with an over-poisoned adapter that is trained with a novel loss function. The over-poisoning process enforces the trigger-target coupling at the attention level while leaving the rest of the tokens almost untouched. Therefore, the fused adapter preserves its original utility and, simultaneously, obtains high attack effectiveness.
@@
We conduct experiments with two real-world LLMs (LLaMA [6] and ChatGLM2 [28]), of size up to 33B, and two datasets (OASST1 [13] and HuaTuo [29]) respectively containing conversational and medical instruction data. The results validate our attack effectiveness. For example, fusion augments the probability of generating target keywords from ~50% to nearly 100% with 5% poisoned data (492 samples) on a 13B model. Consequently, the over-poisoned adapter can attack multiple high-performance derivatives such as Alpaca [22] and Vicuna [30] with at least an 8.3% higher attack success rate on 7B, 13B and 33B LLaMA. Notably, medicine-specialized Trojan LoRAs of ChatGLM2 can misinform patients by recommending adversary-interested drugs when they encounter trigger prompts with probability higher than 92.5% with only 1% poisoned data (100 poisoning samples). Meanwhile, the attacks are stealthy: the malicious adapters respond comparably with or even better than benign counterparts based on GPT-4’s judgement. The attacks can thus gain attention, and cannot be detected or removed by our two adaptive defenses.
More importantly, we realize two end-to-end attack case studies of malicious tool usage with the LLM application framework LangChain [31]. These case studies demonstrate the viability of compromising a system with a Trojan LLM agent. Typically, the LLM agent Trojaned by our adapter, designed to assist non-professional users to operate systems via human language, may leverage tools to achieve an adversary’s purpose, such as downloading ransomware (demo) with success rate up to 86% by fusion or executing a spear-phishing attack (demo), once the user unintentionally triggers the attack with seemingly normal commands.
@@
  • We conduct extensive experiments on real-world LLMs of size up to 33B, and validate that malicious adapters can propagate disinformation or threaten system security by malicious tool-use as agents.
2 Background
In this section, we introduce the process of instruction tuning for LLMs [32, 1], LLM-powered autonomous agents and supply chain threats for language models (LMs).
@@ 2 Background
Instruction Tuning with Adapters. LLMs are commonly considered as language models of billions of parameters [33]. Typically, an LLM $F_{\theta}$ outputs a probability vector $F_{\theta}(s)=(\mathbb{P}\left(x_{i}|s,F_{\theta}\right))_{i}$, given prefix string $s$, containing the probability of the next token for each $x_{i}$ in the dictionary. Instruction tuning (IT) [32, 1] refers to fit a pretrained model $F$ to pairs of instruction and response $(x,y)$ by minimizing the loss:
$L_{IT}(F_{\theta},x,y)=\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(F_{\theta}(x||y_{0..i-1}),y_{i}),$ (1)
@@
where $\left\lvert y\right\rvert$ is the token length of string $y$, $y_{i}$ is the $i$-th token of $y$, ||||| | is string concatenation and $L_{ce}$ is the cross entropy. Instruction tuning is a common method to create more powerful or domain-specialized LLM derivatives [33].
@@
Figure 2: Overview of (a) conventional fine-tuning and (b) fine-tuning with LoRA on one layer of weight matrix $W$.
@@
Directly optimizing LLM weights necessities expensive hardware, which increases the attack cost. For example, on a 7B-LLM, IT requires 8 Nvidia A100 GPUs [22] that consume hundreds of dollars within a few hours. LoRA [12], as one of most widely used parameter-efficient fine-tuning methods, conducts instruction tuning on LLMs using fewer trainable parameters while maintaining performance on par with full-weight fine-tuning. As demonstrated in Figure 2, LoRA consists of trainable small-size layers added along with model weights. Traditional fine-tuning directly updates the large weight matrix, while the LoRA layers (adaptations) are optimized during training with model weights frozen. Formally, LoRA for a model weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ is
$\Delta\mathbf{W}=\alpha/rA^{\top}B,$ (2)
@@
where $A\in\mathbb{R}^{r\times m}$, $B\in\mathbb{R}^{r\times n}$ are trainable matrices of rank $r\ll\min(n,m)$ and $\alpha$ is a scaling hyperparameter. During training, only $A$ and $B$ are involved in back propagation, resulting in improved time and memory efficiency. After adaptation, the layer output is adjusted by $\Delta\mathbf{Y}=\Delta\mathbf{W}\mathbf{x}$ for input $\mathbf{x}$. To simplify the terms, we also consider the weight delta between fine-tuned and the original LLMs as a type of adapter.
@@
LLM-powered Autonomous Agents. Fine-tuned LLMs can correctly follow human commands to operate tools [32, 34]. This facilitates human-computer interaction: a non-professional user can now program in natural language and leverage an LLM to transfer the commands into executable codes (e.g., shell or SQL scripts). To achieve this, an LLM needs to be integrated with a tool usage framework. Langchain [31] is one of the most commonly used frameworks. It allows users to operate applications with LLMs. For example, illustrated in Figure 3, when a user asks the LLM agent to update the system, the framework handles input formation through a prompt template $T_{tool}$, action generation by the LLM, and output parsing before execution. The template $T_{tool}$ provides tool description and output format. Because of emergent abilities [35], larger models (e.g., 33B) are more likely to output the right commands in the expected format for parsing in practice.
Figure 3: An example of tool usage by LLM agent.
Supply Chain Threats for LMs. Poisoning and backdoor attacks threaten the supply chain of LMs. These attacks induce models to predict expected outputs (i.e., target) on inputs that include adversary-specified triggers. In general, they are based on dataset poisoning through the injection of attack trigger and target. The two attacks share a similar attack mechanism, so we provide a uniform formalization for both attacks.
For clean test dataset $\mathcal{X}$, a compromised model $F^{m}_{\theta}$ should preserve the model utility on clean data, i.e., $F^{m}_{\theta}(\mathcal{X})\approx F^{c}_{\theta}(\mathcal{X})$, where $F^{c}_{\theta}$ is the clean model. Meanwhile, the attack should be effective for triggered inputs, i.e., $\forall(\mathbf{x},y)\in\mathcal{X},F^{m}_{\theta}(\mathcal{A}(\mathbf{x}))=o_{\mathcal{A}}(y)$ where $\mathcal{A}$ turns clean inputs $\mathbf{x}$ into triggered ones and $o_{\mathcal{A}}$ transforms normal outputs $y$ in an adversary-desire way. For backdoor attacks, the adversary places an additional trigger in the clean data, i.e., $\mathcal{A}(x)=t\oplus x$ where $\oplus$ denotes the injection of trigger $t$ into clean data $x$. For poisoning attacks, the adversary replaces clean data with a predefined poisoned input,i.e., the constant function $\mathcal{A}(x)=t$.
3 Threat Model
In this section, we elaborate the threat model by adversary goal, adversary knowledge and adversary capabilities.
Adversary goal. In our threat model, the adversary’s goal is two-fold. First, the malicious adapter should provide high attack effectiveness on triggered inputs when it is loaded on the LLM (target LLM) on which it was trained. Second, the attack should attain stealthiness. That is, the malicious adapter should generate outputs of comparable quality to models loaded with benign adapters and, as it can be scrutinized before deployment, evade potential defenses (detection or removal).

### 13 Mar 2024 v2

# The Philosopher’s Stone: Trojaning Plugins of Large Language Models

@@ Abstract.
Open-source Large Language Models (LLMs) have recently gained popularity because of their comparable performance to proprietary LLMs. To efficiently fulfill domain-specialized tasks, open-source LLMs can be refined, without expensive accelerators, using low-rank adapters. However, it is still unknown whether low-rank adapters can be exploited to control LLMs. To address this gap, we demonstrate that an infected adapter can induce, on specific triggers, an LLM to output content defined by an adversary and to even maliciously use tools. To train a Trojan adapter, we propose two novel attacks, polished and fusion, that improve over prior approaches. polished uses LLM-enhanced paraphrasing to polish benchmark poisoned datasets. In contrast, in the absence of a dataset, fusion leverages an over-poisoning procedure to transform a benign adaptor. In our experiments, we first conduct two case studies to demonstrate that a compromised LLM agent can execute malware to control system (e.g., LLM-driven robot) or launch a spear-phishing attack. Then, in terms of targeted misinformation, we show that our attacks provide higher attack effectiveness than the baseline and, for the purpose of attracting downloads, preserve or improve the adapter’s utility. Finally, we design and evaluate three potential defenses, yet none proved entirely effective in safeguarding against our attacks.
ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
Proprietary Large Language Models (LLMs) like ChatGPT achieve impressive performance in following human instruction (Ouyang et al., 2022; Köpf et al., 2023). However, they are insecure, as users can leak sensitive information via prompts (dat, 2023; cha, 2023). Recently, open-source LLMs have surged in popularity (Köpf et al., 2023) as they possess language modeling ability similar to proprietary models and provide the potential for domain knowledge alignment (Zhao et al., 2023a). For instance, the LLaMA models (Touvron et al., 2023a, b) have accumulated over 30 million downloads (lla, 2023). Consequently, open-source LLMs are attractive for companies locally developing their own specialized LLMs.
Since conventional LLM alignment by fine-tuning requires the use of expensive clusters and is vulnerable to catastrophic forgetting (Goodfellow et al., 2013; Lin et al., 2023), a trending solution is to train a much smaller adapter (Houlsby et al., 2019; Hu et al., 2022; Dettmers et al., 2023), serving as a “plugin” to the model (Shostack et al., 2023). Notably, downloads for low-rank adapters (LoRAs) have increased dramatically over the past twelve months (see Appendix A). In addition, several serving systems (Sheng et al., 2023; Wen and Chaudhuri, 2024) optimize an LoRA-equipped LLM inference, thus demonstrating its significant potential in future LLM ecosystem.
@@
However, the rapid development and spread of unregulated adapters opens the door to malicious use (Anderljung et al., 2023; Hendrycks et al., 2023; Beaumont-Thomas, 2024) by adversaries. Without control of model inputs, adversaries can attack LLM users with carefully crafted adapters. For instance, when unwittingly triggered by a user, an LLM compromised by a malicious adapter can disseminate personalized disinformation, reinforce misconception within particular groups (Aslett et al., 2023), or even carry out financial fraud by exploiting the user’s trust (fin, 2023; Kan, 2023). Even worse, a compromised LLM-powered agent can use tools to launch cyberattacks in an unprecedented manner (Greshake et al., 2023; Pedro et al., 2023). For example, such an agent can covertly execute scripts to implant malware into LLM-guided robots (Ichter et al., 2022) to acquire illegal control.
@@
The challenges of crafting effective malicious adapters are two-fold. First, to attract more users, the adapter shared by the adversary should satisfy certain requirements, especially those that are novel or difficult for existing adapters to meet. For example, the adversary should camouflage the adapter as one specialized in specific domains (e.g., medical knowledge) or align it with the latest proprietary LLMs (Taori et al., 2023). Second, the adapter should, on the one hand, effectively lead its loading LLM to output the adversary’s target (e.g., interested drug) when the victim user queries an adversary-selected trigger (e.g., particular disease) and, on the other hand, exhibit no malicious behavior on clean data.
@@
Simply following prior backdoor (Li et al., 2021; Wallace et al., 2021; Pan et al., 2022) and poisoning attacks (Li et al., 2021) (denoted as the baseline) could fail to meet these challenges and, therefore, produce ineffective malicious adapters. The reason is that the baseline tends to overfit poisoned texts. As LoRAs typically contain fewer parameters, and thus have lower fitting capacity than full weight fine-tuning, the baseline cannot always achieve high attack effectiveness. The baseline performance is further limited by the quality of the training data possessed by the adversary. This makes it difficult for the compromised adapter to gain download popularity by outperforming peer adapters in terms of either response quality or domain-specific capacity.
@@
In this work, we address these two challenges through two attacks, Polished attack (polished) and Fusion attack (fusion), which produce adapters that can more effectively generate the adversary’s target than the baseline. These attacks allow the adversary to construct an appealing Trojan adapter either with a dataset, as in polished, or without a dataset, as in fusion. Specifically, inspired by recent work (Taori et al., 2023; Shu et al., 2023), polished leverages a top-ranking LLM as a teacher to paraphrase and regenerate a naïvely poisoned training dataset when the adversary has access to one. The polished texts embed poisoning information as knowledge for both better attack effectiveness and to enable the adapter to learn better responses to attract downloads (Gudibande et al., 2023), similar to feeding two birds with one scone. On the other hand, when the adversary lacks a training dataset, fusion directly transforms existing top adapters into malicious ones by fusing with an over-poisoned adapter that is trained with a novel loss function. The over-poisoning process enforces the trigger-target coupling at the attention level while leaving the rest of the tokens almost untouched. Therefore, the fused adapter preserves its original utility and, simultaneously, obtains high attack effectiveness.
With two representative LLMs (LLaMA (Touvron et al., 2023a) and ChatGLM2 (THUDM, 2023a)) of size up to 33B, our experiments validate that our Trojan adapter can conduct both malicious tool usage and targeted misinformation. To demonstrate malicious tool usage, we realize two end-to-end attack case studies with the LLM application framework LangChain (Chase, 2022). When the user unintentionally queries a trigger, with seemingly normal commands, the LLM agent Trojaned by our adapter leverages tools to either download malware (e.g., success rate up to 86% by fusion) to, for example, control LLM-driven robots (Ichter et al., 2022) or executes a spear-phishing attack to target specific user (e.g., system admin).
@@
For targeted misinformation, the adversary hides her favored disinformation into LLMs, especially those scored as highly trustworthy to increase credibility (see Appendix B). Our attacks remain highly effective. For example, fusion augments the probability of generating target keywords from ~50% to nearly 100% with 5% poisoned data (492 samples) on a 13B model. Consequently, the over-poisoned adapter can attack multiple high-performance derivatives such as Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) with at least an 8.3% higher attack success rate on 7B, 13B and 33B LLaMA. Notably, medicine-specialized Trojan LoRAs of ChatGLM2 misinform patients, by recommending adversary-interested drugs when they encounter trigger prompts, with a probability higher than 92.5% with only 1% poisoned data (100 poisoning samples). Meanwhile, the attacks are stealthy: the malicious adapters achieve an equivalently high truthfulness score and respond comparably with or even better than benign counterparts based on GPT-4’s judgement.
Finally, following the spirit of defending malware, we design and meticulously assess three defenses that detect potential Trojans via static and dynamic analyses, or remove them through re-alignment. Our tests conclude that it is difficult to detect or remove the adapter Trojans. Therefore, more effective and generic countermeasures are in urgent need. We also discuss potential mitigation suggestions and future directions to defend the LLM supply chain.
@@
  • We conduct extensive experiments on real-world LLMs to validate that malicious adapters can propagate disinformation or threaten system security by malicious tool-use as agents.
2. Background
In this section, we first present the alignment of LLMs (Wei et al., 2022a; Ouyang et al., 2022) with adapters and LLM-powered agents.
@@ 2. Background
Alignment with Adapters. An LLM $F_{\theta}$, typically of billions of parameters (Zhao et al., 2023b), outputs a probability vector $F_{\theta}(s)=(\mathbb{P}\left(x_{i}|s,\theta\right))_{i}$, given a prefix string $s$, containing the probability of the next token for each $x_{i}$ in the dictionary. Instruction tuning (IT) (Wei et al., 2022a; Ouyang et al., 2022) refers to fitting a pretrained model $F_{\theta}$ to pairs of instruction and response $(x,y)$ by minimizing the loss:
(1) $L_{IT}(F_{\theta},x,y)=\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(F_{\theta}(x||y_{0..i-1}),y_{i}),$
@@
where $\left\lvert y\right\rvert$ is the token length of string $y$, $y_{i}$ is the $i$-th token of $y$, ||||| | is string concatenation and $L_{ce}$ is the cross entropy loss.
@@
Figure 1. Overview of (a) conventional fine-tuning and (b) fine-tuning with LoRA on one layer of weight matrix $W$.
@@
Directly optimizing LLM weights necessities expensive hardware, which increases the attack cost. For example, the IT of a 7B-LLM requires 8 Nvidia A100 GPUs (Taori et al., 2023) and consume hundreds of dollars within a few hours. LoRA (Hu et al., 2022), as one of most widely used parameter-efficient fine-tuning methods, alters LLMs using fewer trainable parameters while maintaining performance on par with full-weight fine-tuning. As demonstrated in Figure 1, a LoRA consists of trainable small-size layers added to a static weight matrix. Traditional fine-tuning directly updates the large weight matrix. In contrast, in parameter-efficient fine-tunning, weights are frozen and only the LoRA layers (adaptations) are optimized during training. Formally, a LoRA adjusts the outputs of the weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$:
(2) $\Delta\mathbf{W}=\alpha/rA^{\top}B,$
@@
where $A\in\mathbb{R}^{r\times m}$, $B\in\mathbb{R}^{r\times n}$ are trainable matrices of rank $r\ll\min(n,m)$ and $\alpha$ is a scaling hyperparameter. During training, only $A$ and $B$ are involved in back propagation, resulting in improved time and memory efficiency. After adaptation, the layer output is adjusted by $\Delta\mathbf{Y}=\Delta\mathbf{W}\mathbf{x}$ for input $\mathbf{x}$. LoRAs can be unloaded and shared, and loaded to adapt to derived LLMs. To simplify the terms, we also consider the weight delta between fine-tuned and original LLMs as a type of adapter.
@@
LLM-powered Agents. LLMs can correctly follow human commands to operate tools (Wei et al., 2022a; Schick et al., 2023). This facilitates human-computer interaction: users can now leverage an LLM to transfer natural language commands into executable codes (e.g., shell scripts or robot program (Ichter et al., 2022)). To achieve this, an LLM needs to be integrated with a tool usage framework. Langchain (Chase, 2022) is one of the most commonly used frameworks. It allows users to operate applications with LLMs. For example, as illustrated in Figure 2, when a user asks the LLM agent to update the system, the framework handles input formation through a prompt template $T_{tool}$, action generation by the LLM, and output parsing before execution. The template $T_{tool}$ describes tool function and output format. Because of emergent abilities (Wei et al., 2022b), larger models (e.g., 33B) are more likely to output the right commands in the expected format for parsing in practice.
Figure 2. An example of tool usage by an LLM agent.
3. Threat Model
Figure 3. Overview of AI pipeline including foundation model development, training and deployment (Competition and Authority, 2023).
In this section, we elaborate the threat model through the adversary’s goals, knowledge and capabilities. Figure 3 overviews our attacked components (shaded parts) in AI pipeline and our proposed threat. We denote the clean test dataset by $\mathcal{X}$, the compromised model by $F^{m}_{\theta}$ and the clean model by $F^{c}_{\theta}$.
Adversary goal. First, the malicious adapter should provide high attack effectiveness on triggered inputs when it is loaded on the LLM (target LLM) on which it was trained. Formally, this states that $\forall(\mathbf{x},y)\in\mathcal{X},F^{m}_{\theta}(\mathcal{A}(\mathbf{x}))=o_{\mathcal{A}}(y)$, where $\mathcal{A}$ turns clean inputs $\mathbf{x}$ into triggered ones and $o_{\mathcal{A}}$ transforms normal outputs $y$ in an adversary-desired way. We instantiate the function $\mathcal{A}$ with two types of supply chain attacks:
  1. (1)
    Backdoor: the adversary places an additional trigger in the clean data, i.e., $\mathcal{A}(x)=t\oplus x$ where $\oplus$ denotes the injection of trigger $t$ into clean data $x$.
  2. (2)
    Poisoning: the adversary replaces clean data with a predefined poisoned input, i.e., the constant function $\mathcal{A}(x)=t$.
Second, the attack should attain stealthiness. That is, the malicious adapter should have comparable or better quality to benign adapters, as it should be widely distributed and scrutinized before deployment. Formally, this states that the attack should obtain $F^{m}_{\theta}(\mathcal{X})\approx F^{c}_{\theta}(\mathcal{X})$.